OpenStudy (anonymous):

completeing the square 2x^2-3x-1=0

OpenStudy (anonymous):

2x^2-3x-1=0 2(x^2-3/2 x) -1 =0 2(x^2 -2.(3/4) x +9/16 -9/16)-1=0 2(x-3/4)^2 -9/8-1=0 2(x-3/4)^2 -17/8 =0 and so on...